How do I find the slope of the equation 10x - 21y = 0? I first tried to put it in the form of y = mx + b, resulting in y = 10x + 21. Afterwards I put 10/21, but there is no way to reduce and 10/21 isn't an option for my answer.
Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

How do I find the slope of the equation 10x - 21y = 0? I first tried to put it in the form of y = mx + b, resulting in y = 10x + 21. Afterwards I put 10/21, but there is no way to reduce and 10/21 isn't an option for my answer.
-----------------
10x - 21y = 0?
y = (10/21)x [not 10x + 21]
Slope = 10/21
That's the slope, regardless of the options.

Answer by rothauserc(4718)

We are given
How do I find the slope of the equation 10x - 21y = 0?
first solve the equation for y
21y = 10x
and
y = (10/21) * x
now you can use y = mx + b and we have
m = 10 / 21